# Magnetic field and kinetic helicity evolution in simulations of interacting disk galaxies

### v1

@@ Abstract
Aims. We carry out a parameter study of interacting disk galaxies with impact parameters ranging from central collisions to weakly interacting scenarios. The orientations of the disks are also varied. In particular, we investigate how magnetic field amplification depends on the these parameters.
Methods. We use magnetohydrodynamics for gas disks in combination with live dark matter halos in adaptive mesh refinement simulations. The disks are initialized using a setup for isolated disks in hydrostatic equilibrium. Since we focus on the impact of tidal forces on magnetic field evolution, adiabatic physics is applied. Small-scale filtering of the velocity and magnetic field allows us to estimate the turbulent electromotive force (EMF) and kinetic helicity.
Results. Time series of the average magnetic field in central and outer disk regions show pronounced peaks during close encounters and mergers. This agrees with observed magnetic fields at different interaction stages. The central field strength exceeds $10\;\upmu$G (corresponding to an amplification factor of 2 to 3) for small impact parameters. As the disks are increasingly disrupted and turbulence is produced by tidal forces, the small-scale EMF reaches a significant fraction of the total EMF. The small-scale kinetic helicity is initially antisymmetric across the disk plane. Albeit its evolution is sensitive to both the impact parameter and inclinations of the rotation axes with respect to the relative motion of the disks, antisymmetry is generally broken through interactions and the merger remnant has lost most of the initial helicity. The EMF and the magnetic field also decay rapidly after coalescence.
Conclusions. The strong amplification during close encounters of the interacting galaxies are mostly driven by helical flows and a mean-field dynamo. The small-scale dynamo contributes significantly in post-interaction phases. However, the amplification of the magnetic field cannot be sustained.
@@ 1 Introduction
Observations show that galaxy interactions and mergers play a significant role during the evolution of galaxies (Struck 1999; Smith et al. 2010). In addition to an enhanced star formation rate (Kennicutt 1998; Georgakakis et al. 2000; Rossa et al. 2007), there are indications that mergers have an impact on the magnetic field strength in galaxies (Condon et al. 2002; Chyży & Beck 2004; Drzazga et al. 2011), suggesting an amplification of the field due to dynamo action. While the $\alpha\Omega$-dynamo induced by the differential rotation of disk galaxies is thought to play an important role for the generation of galactic magnetic fields from a weak primordial field (Brandenburg & Subramanian 2005; Park et al. 2013; Steinwandel et al. 2020; Wissing & Shen 2023), the $\alpha^{2}$ dynamo caused by inhomogeneous shear flows with kinetic helicity and the turbulent small-scale dynamo induced by non-helical, isotropic turbulence on small scales are expected to contribute significantly to the field amplification in merging galaxies (Arshakian et al. 2009; Schober et al. 2013; Bhat & Subramanian 2014; Moss et al. 2014). In contrast to isolated disk galaxies, where turbulence is assumed to be driven by supernova explosions (Hopkins et al. 2013) or gravitational instabilities (Krumholz & Burkhart 2016), shear and tidal effects can produce additional turbulence during galaxy interactions (Renaud et al. 2014; Whittingham et al. 2021; Jiménez et al. 2023).
@@
The interaction process has been extensively studied in idealized numerical simulations without cosmological initial conditions. The basic idea is to embed disks composed of gas and possibly stars in spherical dark matter halos and to set them on a collision course. An original study of such systems using a simple particle-based model was presented by Toomre & Toomre (1972) five decades ago. More recently, hybrid N-body/smoothed particle hydrodynamics (SPH) simulations Mihos & Hernquist (1996); Di Matteo et al. (2008); Blumenthal & Barnes (2018) and semi-Lagrangian simulations (Hayward et al. 2014; Moreno et al. 2019) focused on the dynamics of the interstellar medium (ISM) and interaction-induced star formation. Merger simulations with magnetic fields were first performed by Kotarba et al. (2010), using a magnetohydrodynamics (MHD) extension of SPH (Dolag & Stasyszyn 2009). Their simulation of the Antennae galaxies suggests that mergers play a significant role in the amplification of galactic magnetic fields. Magnetic field amplification during minor mergers, where the secondary galaxy has only a small fraction of the mass of the primary galaxy, was investigated with a similar approach by Geng et al. (2012) and in cosmological simulations with the semi-Lagrangian Arepo code by Pakmor et al. (2014). Morphological changes due to mergers were studied for a sample of galaxies from the Horizon-AGN cosmological simulation (Martin et al. 2018). Moreover, the role of magnetic fields and the small-scale dynamo was analysed in great detail in zoom-in simulations of major mergers from the Illustris simulation by Whittingham et al. (2021). However, in contrast to simulating pairs of galaxies, the resolution of cosmological simulations is limited and interaction parameters cannot be varied systematically.
@@
While Lagrangian methods undoubtedly have merits for the treatment of moving galaxies, Rodenbeck & Schleicher (2016) demonstrated that merger simulations are feasible in a greatly simplified setup with purely gaseous and adiabatic disks without dark matter halos using the magnetohydrodynamical adaptive mesh refinement (AMR) code Enzo (Bryan et al. 2014). They applied the method of Wang et al. (2010) to initialize an isolated equilibrium disk in a static external potential representing the gravity of dark matter with additional support by magnetic pressure. However, they excluded the external potential because it is not applicable to moving disks in a galaxy merger scenario, although the contribution of dark matter is dominant. Even so, this approach avoids the difficulties of implementing MHD into particle-based codes (Dolag & Stasyszyn 2009) and, owing to a well defined resolution scale, offers the advantage of better controllability of the small-scale physics. For realistic simulations of interacting galaxies, however, N-body dynamics is essential. Hydrodynamical AMR simulations of a major merger similar to the Antennae system with live N-body halos were performed by Teyssier et al. (2010) and Renaud et al. (2015). At spatial resolutions in the parsec range, they were able to analyse the production of turbulence in the ISM due to tidal interactions and the compression-induced enhancement of star formation.
@@
In this paper, we combine the approaches of Teyssier et al. (2010) and Rodenbeck & Schleicher (2016). To generate initial conditions for spherical dark matter halos, we follow Drakos et al. (2017). The gaseous disks are based on Wang et al. (2010). In essence, the gas density and rotation velocity of a disk in hydrostatic equilibrium are determined iteratively from the self-gravity of the gas and the smoothed potential of the dark-matter halo. The advantage of this method of initializing disks is that effects caused by the initial relaxation are reduced and changes in the disk structure are mainly due to physical effects, such as tidal forces and fragmentation induced by cooling and gravity. The applied disk and halo models are described in more detail in the following section. In Sect. 3, the numerical setup is outlined and an overview of our simulation suite is provided. In this study, we focus on magnetic field amplification due to tidal interactions between the disks and their halos. While radiative cooling and stellar feed-back are essential ingredients for a realistic disk structure, the findings of Ntormousi et al. (2020) indicate that these processes are of minor importance for the generation of galactic magnetic fields. Since non-adiabatic physics increases computational requirements substantially for the targeted spatial resolution of about 101010\;10pc, we use adiabatic gas dynamics in this work. The analysis of the simulation is presented in Sect. 4: After a phenomenological discussion of representative examples, the average magnetic field in central and outer disk regions are compared among different simulations and subsequent encounters of the interacting galaxies in each simulation. Moreover, the question whether magnetic field amplification during interaction phases can be related dynamo action on the smallest numerically resolved length scales is investigated. In the conclusions, we sum up and discuss our results.
@@ 2 Disk and halo models
To initialize the gaseous disks, we apply the potential method of Wang et al. (2010) for a disk with exponential surface density:
@@
where $r$ and $z$ are, respectively, radius and height in a cylindrical coordinate system with the disk midplane at $z=0$. The surface density at the center is given by $\Sigma_{0}=M_{\mathrm{d}}/2\uppi r_{\mathrm{d}}^{2}$ for a disk of total mass $M_{\mathrm{d}}$ and scale length $r_{\mathrm{d}}$. In many isolated disk simulations, an exponential form is assumed for the volume density $\rho(r,z)$ of the gas. However, as argued by Wang et al. (2010), this implies a ring in the surface density.
@@
$c_{0}^{2}:=\frac{p}{\rho}=\frac{k_{\mathrm{B}}T}{\mu m_{\mathrm{H}}}$ (3)
@@
Equation (1) is satisfied if the midplance density $\rho_{0}(r)=\rho(r,0)$ is given by
$\rho_{0}(r)=\frac{\Sigma(r)}{\int_{\infty}^{\infty}\exp{\left(-\Phi_{z}(r,z)/[(1+\epsilon_{\mathrm{m}})c_{0}^{2}]\right)}\mathrm{d}z}\,,$ (7)
The integral in the denominator can be interpreted as a disk scale height that varies with radius and is determined by the model. Equations (6) and (7) cannot be solved analytically. A numerical solution is readily obtained by using an iterative scheme in combination with a numerical solver for the differential equation (see Sect. 3). For an overview of all model parameters, see Table 1.
@@
For the dark matter halo, we assume the radial density profile (Hernquist 1990)
@@
where $a$ is the scale length of the halo and $M_{\mathrm{dm}}$ the total cumulative mass for $r\rightarrow\infty$. Here, $r$ is a spherical radial coordinate. As proposed by Springel et al. (2005), the scale length of the Hernquist halo is chosen such that it contains the same mass $M_{200}$ within the virial radius $r_{200}$ as an NFW halo of given scale length $r_{\mathrm{s}}$ and concentration parameter c=$r_{200}/r_{\mathrm{s}}$:
@@
We use the method developed by Drakos et al. (2017) to compute random initial positions and velocities of the particles. The algorithm populates the phase space of an N-body system based on the distribution function of a halo in virial equilibrium.
@@
$v_{\mathrm{dm}}(r)=\frac{\sqrt{GM_{\mathrm{dm}}r}}{r+a}\,.$ (12)
@@
and $B_{r}=B_{z}=0$ in the disk’s cylindrical coordinate system. The disk interior is defined by $\rho\geq\rho_{\mathrm{min}}$, where $\rho_{\mathrm{min}}=6.768\times 10^{-30}\;\mathrm{g}\;\mathrm{cm}^{-3}$ is the constant density of the ambient medium. While we accounted for magnetic pressure in the computation of the equilibrium density and rotation curve, a toroidal magnetic field gives rise to an additional curvature force that is directed radially inwards. In the case of a constant plasma beta, this force gives rise to a constant negative contribution to the rotation curve. We neglected this contribution because it is lower by several orders of magnitude than the maximum rotation speed for the chosen disk parameters.
@@
  • The density of the gaseous disk given by eq. (4) decreases rapidly with the scale height. To avoid unphysically low densities and density jumps at the interface between disk and ambient medium, the density floor $\rho_{\mathrm{min}}$ mentioned above is applied when setting up the disk. Since a constant density is not consistent with the equilibrium solution, gas is subsequently falling toward the disk. This could be alleviated by modelling a circumgalactic medium (see, for example, Steinwandel et al. 2019). However, the total mass in the ambient medium is very small compared to the disk mass.
@@
As a consequence, perturbations of the initial equilibrium disk are unavoidable. For the most part, the disk does not undergo abrupt changes, but gradually (over a few 100100100\;100Myr) evolves into a disk that exhibits a spiral-like structure.
@@ 3 Numerical methods and overview of simulations
We use a modified version of the publicly available cosmological AMR code Enzo (Bryan et al. 2014; Brummel-Smith et al. 2019). [^1] See the website enzo-project.org for further details. For the simulation suite presented in this article, the public repository github.com/SimonCSelg/enzo-dev_mhdsgs-diskgalaxy, commit fc68caa, was used. In order to maintain consistency between the simulations, this repository was frozen. Subsequent modifications are publicly available at github.com/wolfram-schmidt/enzo-dev/tree/diskgalaxy. The code is MPI parallelized, features N-body dynamics based on a second-order drift-kick-drift algorithm in combination with cloud-in-cell interpolation to compute the joint gravitational potential of gas and particles, and a variety of different finite volume solvers for gas dynamics and magnetic fields. In our simulations, we apply the monotonic upstream-centered scheme for conservation laws (MUSCL) with a local Lax-Friedrichs (LLF) solver for sufficient numerical robustness. Adiabatic gas dynamics is applied under the assumption that magnetic field dynamics during interactions is mainly driven by tidal forces. Since radiative cooling, star formation, and feedback processes are not included, the spatial resolution and time stepping requirements are significantly reduced. Even so, the required computational resources for an extensive parameter study were substantial, as detailed below.
@@
Originally, the computation of initial conditions for a disk in hydrostatic equilibirum were part of the initialization in Enzo (Rodenbeck & Schleicher 2016). However, this was not practical for the particle initial conditions required for the live dark-matter halo. For this reason, we implemented a Python tool for computing particle initial conditions along the lines of ICICLE (Drakos et al. 2017). By applying object-oriented programming, we were readily able to incorporate equilibrium disks as a subclass of the halo class into our Python code. The algorithms outlined in Sect. 2 are implemented as object methods and make use of the highly optimized numerical methods for root finding and integration from the SciPy library. The resulting disk and particle data can be saved to files, which in turn can be read by Enzo to set up one or more disk galaxies embedded in live dark-matter halos. [^2] The Python modules for halos and equilibrium disks are publicly available at github.com/wolfram-schmidt/isolated-galaxy.
@@
We restrict ourselves to interactions of Milkyway-like galaxies of equal mass. The parameters chosen for our simulations are summarized in Table 1. Each galaxy initially consists of a gas disk of $10^{10}\;\mathrm{M}_{\odot}$ and a dark matter halo of $M_{200}=10^{12}\;M_{\odot}$, where $M_{200}$ is the mass enclosed within the virial radius $r_{200}=210\;$kpc. The radial scale length of the disks is $r_{\mathrm{d}}=3.5\;$kpc. The intial conditions were computed using the models outlined in Section 2. The initial separation $d_{\mathrm{sep},0}$ of the two galaxies is set to the cut-off radius $r_{200}$ of a dark matter halo. For a periodic box of 222\;2Mpc linear size, this choice ensures that the separation of the centers of mass is small compared to the distance from the periodic boundaries. Although the halos are overlapping, the initial separation of the gas disks is significantly larger than their size. For our production runs, we use a $512^{3}$ root grid and 8 levels of refinement by overdensity and shear, corresponding to a maximal spatial resolution of 151515\;15pc. The application of the second refinement criterion is important to cover interaction zones with high resolution, especially between the galaxies and in the outer disk regions where the gas density is relatively low. However, this greatly increases the computational cost of the simulations. [^3] Using about 1200 MPI tasks, typically 111\;1million core-h were needed to complete one simulation. Owing to inherent limitations of the parallelization scheme used in Enzo, it is not feasible to run the code with more than a few thousand tasks.
@@
  • the initial translational velocities ${V}_{1}$ and ${V}_{2}$ of the two disks,
  • the impact parameter $b$, i. e., the normal distance between a straight line in the direction of the initial velocity of the secondary galaxy, ${V}_{2}$, and the center of the primary galaxy,
@@
In our simulation suite, we examine a series of impact parameters combined with varying orientations of the galaxies. To specify the orientations of the galaxies, we use the inclination angles of each galaxy’s angular momentum vector with respect to the vector of the initial relative velocity of the interacting system, ${V}_{\mathrm{rel}}=\@vec{V}_{1}-\@vec{V}_{2}$:
@@
Table 2: Overview of varied simulation parameters and minimum distances between galaxy centers at the first, second, and third encounter, respectively. Dashes correspond to missing data in the case of simulations that were not evolved long enough to undergo more than one or two encounters. The parameters are the inclination angles $i_{1}$ and $i_{2}$ of the disk’s rotation axes measured with respect to their initial relative velocity ${V}_{\mathrm{rel}}$, and the angle $\alpha_{\mathrm{b}}$ defined by $\sin{\alpha_{\mathrm{b}}}=b/d_{\mathrm{sep},0}$, where $b$ is the impact parameter and $d_{\mathrm{sep},0}$ the initial separation of the two disks.
@@
    Group III: face-on collisions with relative velocity parallel/anti-parallel to rotation axes
@@
In each group, the impact parameter $b$ is varied between the limiting scenarios of a central collision ($b=0$) and an interaction where the galaxies and their halos are moderately distorted by tidal forces ($b$ of the order of virial radius). In the following, we use the angle defined by $\sin{\alpha_{\mathrm{b}}}=b/d_{\mathrm{sep,0}}$ to specify the impact parameter. For the absolute value of the initial relative velocity, we assume |${V}_{\mathrm{rel}}|=v_{200}$ in all cases.
Owing to the required computational time and substantial storage requirements, only a few dozens scenarios with more or less arbitrary parameter choices were feasible. Thus, our simulation suite covers some typical scenarios, such as edge-on (group 0) and face-on (group III) collisions of aligned disks, and a selection of intermediate cases with non-zero relative inclinations of the disks. In groups IV and V we investigate scenarios where the secondary galaxy moves along an inbound trajectory that is inclined with respect to the disk plane of the primary (similar to satellites of the Milky Way, such as the Magellanic Clouds). If feasible, the simulations were run until the post-merger stage. In group 0, this was achieved for Sim. No. 1 to 5 within 2.52.52.5\;2.5Gyr, but not for Sim. No. 6. In the other groups, reductions of the fraction of the time step implied by the CFL criterion required terminating the simulations already after 1.51.51.5\;1.5Gyr in order to save computational time. However, all simulations captured at least the first interaction between the galaxies.

### v2

@@ Abstract
Aims. We carried out a parameter study of interacting disk galaxies with impact parameters ranging from central collisions to weakly interacting scenarios. The orientations of the disks were also varied. In particular, we investigated how magnetic field amplification depends on these parameters.
Methods. We used magnetohydrodynamics for gas disks in combination with live dark-matter halos in adaptive mesh refinement simulations. The disks were initialized using a setup for isolated disks in hydrostatic equilibrium. Since we focused on the impact of tidal forces on magnetic field evolution, adiabatic physics was applied. Small-scale filtering of the velocity and magnetic field allowed us to estimate the turbulent electromotive force (EMF) and kinetic helicity.
Results. Time series of the average magnetic field in central and outer disk regions show pronounced peaks during close encounters and mergers. This agrees with observed magnetic fields at different interaction stages. The central field strength exceeds $10\;\upmu$G (corresponding to an amplification factor of 2 to 3) for small impact parameters. As the disks are increasingly disrupted and turbulence is produced by tidal forces, the small-scale EMF reaches a significant fraction of the total EMF. The small-scale kinetic helicity is initially antisymmetric across the disk plane. Though its evolution is sensitive to both the impact parameter and inclinations of the rotation axes with respect to the relative motion of the disks, antisymmetry is generally broken through interactions and the merger remnant loses most of the initial helicity. The EMF and the magnetic field also decay rapidly after coalescence.
Conclusions. The strong amplification during close encounters of the interacting galaxies is mostly driven by helical flows and a mean-field dynamo. The small-scale dynamo contributes significantly in post-interaction phases. However, the amplification of the magnetic field cannot be sustained.
@@ 1 Introduction
Observations show that galaxy interactions and mergers play a significant role during the evolution of galaxies (Struck 1999; Smith et al. 2010). In addition to increasing the star formation rate (Kennicutt 1998; Georgakakis et al. 2000; Rossa et al. 2007), there are indications that mergers have an impact on the magnetic field strength in galaxies (Condon et al. 2002; Chyży & Beck 2004; Drzazga et al. 2011), suggesting an amplification of the field due to dynamo action. While the $\alpha$-$\Omega$ dynamo induced by the differential rotation of disk galaxies is thought to play an important role in the generation of galactic magnetic fields from a weak primordial field (Brandenburg & Subramanian 2005; Park et al. 2013; Steinwandel et al. 2020; Wissing & Shen 2023), the $\alpha^{2}$ dynamo (caused by inhomogeneous shear flows with kinetic helicity) and the turbulent small-scale dynamo (induced by non-helical, isotropic turbulence on small scales) are expected to contribute significantly to the field amplification in merging galaxies (Arshakian et al. 2009; Schober et al. 2013; Bhat & Subramanian 2014; Moss et al. 2014). In contrast to isolated disk galaxies, where turbulence is assumed to be driven by supernova explosions (Hopkins et al. 2013) or gravitational instabilities (Krumholz & Burkhart 2016), shear and tidal effects can produce additional turbulence during galaxy interactions (Renaud et al. 2014; Whittingham et al. 2021; Jiménez et al. 2023).
@@
The interaction process has been extensively studied in idealized numerical simulations without cosmological initial conditions. The basic idea is to embed disks composed of gas and possibly stars in spherical dark-matter halos and to set them on a collision course. An original study of such systems using a simple particle-based model was presented by Toomre & Toomre (1972) five decades ago. More recently, smoothed particle hydrodynamic simulations (Mihos & Hernquist 1996; Di Matteo et al. 2008; Blumenthal & Barnes 2018) and semi-Lagrangian simulations (Hayward et al. 2014; Moreno et al. 2019) focused on the dynamics of the interstellar medium and interaction-induced star formation. Merger simulations with magnetic fields were first performed by Kotarba et al. (2010), using a magnetohydrodynamic (MHD) extension of smoothed particle hydrodynamics (Dolag & Stasyszyn 2009). Their simulation of the Antennae galaxies suggests that mergers play a significant role in the amplification of galactic magnetic fields. Magnetic field amplification during minor mergers, where the secondary galaxy has only a small fraction of the mass of the primary galaxy, was investigated with a similar approach by Geng et al. (2012) and in cosmological simulations with the semi-Lagrangian Arepo code by Pakmor et al. (2014). Morphological changes due to mergers were studied for a sample of galaxies from the Horizon-AGN cosmological simulation (Martin et al. 2018). Moreover, the role of magnetic fields and the small-scale dynamo was analyzed in great detail in zoom-in simulations of major mergers from the Illustris simulation by Whittingham et al. (2021). However, in contrast to simulating pairs of galaxies, the resolution of cosmological simulations is limited and interaction parameters cannot be varied systematically.
@@
While Lagrangian methods undoubtedly have merits for the treatment of moving galaxies, Rodenbeck & Schleicher (2016) demonstrate that merger simulations are feasible in a greatly simplified setup with purely gaseous and adiabatic disks without dark-matter halos using the MHD adaptive mesh refinement (AMR) code Enzo (Bryan et al. 2014). They applied the method from Wang et al. (2010) to initialize an isolated equilibrium disk in a static external potential that represents the gravity of dark matter with additional support from magnetic pressure. However, they excluded the external potential because it is not applicable to moving disks in a galaxy merger scenario, although the contribution of dark matter is dominant. Even so, this approach avoids the difficulties of implementing MHD into particle-based codes (Dolag & Stasyszyn 2009) and, owing to a well-defined resolution scale, offers the advantage of better controllability of the small-scale physics. For realistic simulations of interacting galaxies, however, N-body dynamics is essential. Hydrodynamic AMR simulations of a major merger similar to the Antennae system with live N-body halos were performed by Teyssier et al. (2010) and Renaud et al. (2015). At spatial resolutions in the parsec range, they were able to analyze the production of turbulence in the interstellar medium due to tidal interactions and the compression-induced enhancement of star formation.
@@
In this paper, we combine the approaches of Teyssier et al. (2010) and Rodenbeck & Schleicher (2016). To generate initial conditions for spherical dark-matter halos, we follow Drakos et al. (2017). The gaseous disks are based on Wang et al. (2010). In essence, the gas density and rotation velocity of a disk in hydrostatic equilibrium are determined iteratively from the self-gravity of the gas and the smoothed potential of the dark-matter halo. The advantage of this method of initializing disks is that the effects of the initial relaxation are reduced and changes in the disk structure are mainly due to physical effects, such as tidal forces and fragmentation induced by cooling and gravity. The applied disk and halo models are described in more detail in the following section. In Sect. 3, the numerical setup is outlined and an overview of our simulation suite is provided. In this study we focus on magnetic field amplification due to tidal interactions between the disks and their halos. While radiative cooling and stellar feedback are essential ingredients for a realistic disk structure, the findings of Ntormousi et al. (2020) indicate that these processes are of minor importance for the generation of galactic magnetic fields. Since nonadiabatic physics increases computational requirements substantially for the targeted spatial resolution of about 101010\;10pc, we use adiabatic gas dynamics in this work. The analysis of the simulation is presented in Sect. 4: After a phenomenological discussion of representative examples, the average magnetic fields in central and outer disk regions from different simulations and subsequent encounters of the interacting galaxies in each simulation are compared. Moreover, the question of whether magnetic field amplification during interaction phases can be related to dynamo action on the smallest numerically resolved length scales is investigated. In the conclusions, we sum up and discuss our results.
@@ 2 Disk and halo models
To initialize the gaseous disks, we applied the potential method of Wang et al. (2010) for a disk with exponential surface density:
@@
where $r$ and $z$ are, respectively, the radius and height in a cylindrical coordinate system with the disk midplane at $z=0$. The surface density at the center is given by $\Sigma_{0}=M_{\mathrm{d}}/2\uppi r_{\mathrm{d}}^{2}$ for a disk of total mass $M_{\mathrm{d}}$ and scale length $r_{\mathrm{d}}$. In many isolated disk simulations, an exponential form is assumed for the volume density $\rho(r,z)$ of the gas. However, as argued by Wang et al. (2010), this implies a ring in the surface density.
@@
$c_{0}^{2}:=\frac{p}{\rho}=\frac{k_{\mathrm{B}}T}{\mu m_{\mathrm{H}}},$ (3)
@@
Equation (1) is satisfied if the midplane density $\rho_{0}(r)=\rho(r,0)$ is given by
$\rho_{0}(r)=\frac{\Sigma(r)}{\int_{\infty}^{\infty}\exp{\left(-\Phi_{z}(r,z)/[(1+\epsilon_{\mathrm{m}})c_{0}^{2}]\right)}\mathrm{d}z}\,.$ (7)
The integral in the denominator can be interpreted as a disk scale height that varies with radius and is determined by the model. Equations (6) and (7) cannot be solved analytically. A numerical solution is readily obtained by using an iterative scheme in combination with a numerical solver for the differential equation (see Sect. 3). An overview of all model parameters is provided in Table 1.
@@
For the dark-matter halo, we assumed the radial density profile (Hernquist 1990)
@@
where $a$ is the scale length of the halo and $M_{\mathrm{dm}}$ the total cumulative mass for $r\rightarrow\infty$. Here, $r$ is a spherical radial coordinate. As proposed by Springel et al. (2005), the scale length of the Hernquist halo is chosen such that it contains the same mass $M_{200}$ within the virial radius $r_{200}$ as a Navarro–Frenk–White halo of given scale length $r_{\mathrm{s}}$ and concentration parameter c=$r_{200}/r_{\mathrm{s}}$:
@@
We used the method developed by Drakos et al. (2017) to compute random initial positions and velocities of the particles. The algorithm populates the phase space of an N-body system based on the distribution function of a halo in virial equilibrium.
@@
$v_{\mathrm{dm}}(r)=\frac{\sqrt{GM_{\mathrm{dm}}r}}{r+a}\,$ (12)
@@
and $B_{r}=B_{z}=0$ in the disk’s cylindrical coordinate system. The disk interior is defined by $\rho\geq\rho_{\mathrm{min}}$, where $\rho_{\mathrm{min}}=6.768\times 10^{-30}\;\mathrm{g}\;\mathrm{cm}^{-3}$ is the constant density of the ambient medium. While we accounted for magnetic pressure in the computation of the equilibrium density and rotation curve, a toroidal magnetic field gives rise to an additional curvature force that is directed radially inward. In the case of a constant plasma beta, this force gives rise to a constant negative contribution to the rotation curve. We neglected this contribution because it is lower by several orders of magnitude than the maximum rotation speed for the chosen disk parameters.
@@
  • The density of the gaseous disk given by Eq. (4) decreases rapidly with the scale height. To avoid unphysically low densities and density jumps at the interface between disk and ambient medium, the density floor $\rho_{\mathrm{min}}$ mentioned above is applied when setting up the disk. Since a constant density is not consistent with the equilibrium solution, gas is subsequently falling toward the disk. This could be alleviated by modeling a circumgalactic medium (see, for example, Steinwandel et al. 2019). However, the total mass in the ambient medium is very small compared to the disk mass.
@@
As a consequence, perturbations of the initial equilibrium disk are unavoidable. For the most part, the disk does not undergo abrupt changes, but gradually (over a few hundred megayears) evolves into a disk that exhibits a spiral-like structure.
@@ 3 Numerical methods and overview of simulations
We used a modified version of the publicly available cosmological AMR code Enzo (Bryan et al. 2014; Brummel-Smith et al. 2019). [^1] See the website enzo-project.org for further details. For the simulation suite presented in this article, the public repository github.com/SimonCSelg/enzo-dev_mhdsgs-diskgalaxy, commit fc68caa, was used. In order to maintain consistency between the simulations, this repository was frozen. Subsequent modifications are publicly available at github.com/wolfram-schmidt/enzo-dev/tree/diskgalaxy. The code is MPI-parallelized, features N-body dynamics based on a second-order drift-kick-drift algorithm in combination with cloud-in-cell interpolation to compute the joint gravitational potential of gas and particles, and a variety of different finite volume solvers for gas dynamics and magnetic fields. In our simulations, we applied the monotonic upstream-centered scheme for conservation laws with a local Lax-Friedrichs solver for sufficient numerical robustness. Adiabatic gas dynamics is applied under the assumption that magnetic field dynamics during interactions is mainly driven by tidal forces. Since radiative cooling, star formation, and feedback processes are not included, the spatial resolution and time stepping requirements are significantly reduced. Even so, the required computational resources for an extensive parameter study were substantial, as detailed below.
@@
Originally, the computation of initial conditions for a disk in hydrostatic equilibrium were part of the initialization in Enzo (Rodenbeck & Schleicher 2016). However, this was not practical for the particle initial conditions required for the live dark-matter halo. For this reason, we implemented a Python tool for computing particle initial conditions along the lines of ICICLE (Drakos et al. 2017). By applying object-oriented programming, we were readily able to incorporate equilibrium disks as a subclass of the halo class into our Python code. The algorithms outlined in Sect. 2 are implemented as object methods and make use of the highly optimized numerical methods for root finding and integration from the SciPy library. The resulting disk and particle data can be saved to files, which in turn can be read by Enzo to set up one or more disk galaxies embedded in live dark-matter halos. [^2] The Python modules for halos and equilibrium disks are publicly available at github.com/wolfram-schmidt/isolated-galaxy.
@@
We restricted ourselves to interactions of Milky Way-like galaxies of equal mass. The parameters chosen for our simulations are summarized in Table 1. Each galaxy initially consists of a gas disk of $10^{10}\;\mathrm{M}_{\odot}$ and a dark-matter halo of $M_{200}=10^{12}\;M_{\odot}$, where $M_{200}$ is the mass enclosed within the virial radius $r_{200}=210\;$kpc. The radial scale length of the disks is $r_{\mathrm{d}}=3.5\;$kpc. The intial conditions were computed using the models outlined in Section 2. The initial separation $d_{\mathrm{sep},0}$ of the two galaxies is set to the cut-off radius $r_{200}$ of a dark-matter halo. For a periodic box of 222\;2Mpc linear size, this choice ensures that the separation of the centers of mass is small compared to the distance from the periodic boundaries. Although the halos are overlapping, the initial separation of the gas disks is significantly larger than their size. For our production runs, we used a $512^{3}$ root grid and 8 levels of refinement by overdensity and shear, corresponding to a maximal spatial resolution of 151515\;15pc. The application of the second refinement criterion is important to cover interaction zones with high resolution, especially between the galaxies and in the outer disk regions where the gas density is relatively low. However, this greatly increases the computational cost of the simulations. [^3] Using about 1200 MPI tasks, typically 111\;1million core-h were needed to complete one simulation. Owing to inherent limitations of the parallelization scheme used in Enzo, it is not feasible to run the code with more than a few thousand tasks.
@@
  • the initial translational velocities of the two disks, ${V}_{1}$ and ${V}_{2}$ ,
  • the impact parameter, $b$, which is the normal distance between a straight line in the direction of the initial velocity of the secondary galaxy, ${V}_{2}$, and the center of the primary galaxy,
@@
In our simulation suite, we examined a series of impact parameters combined with varying orientations of the galaxies. To specify the orientations of the galaxies, we used the inclination angles of each galaxy’s angular momentum vector with respect to the vector of the initial relative velocity of the interacting system, ${V}_{\mathrm{rel}}=\@vec{V}_{1}-\@vec{V}_{2}$:
@@
Table 2: Overview of varied simulation parameters and minimum distances between galaxy centers.
@@
 [^4] The parameters are the inclination angles $i_{1}$ and $i_{2}$ of the disk’s rotation axes measured with respect to their initial relative velocity ${V}_{\mathrm{rel}}$, and the angle $\alpha_{\mathrm{b}}$ defined by $\sin{\alpha_{\mathrm{b}}}=b/d_{\mathrm{sep},0}$, where $b$ is the impact parameter and $d_{\mathrm{sep},0}$ the initial separation of the two disks. Dashes correspond to missing data in the case of simulations that were not evolved long enough to undergo more than one or two encounters.
@@
    Group III: face-on collisions with relative velocity parallel or antiparallel to rotation axes
@@
In each group, the impact parameter $b$ is varied between the limiting scenarios of a central collision ($b=0$) and an interaction where the galaxies and their halos are moderately distorted by tidal forces ($b$ on the order of virial radius). In the following, we use the angle defined by $\sin{\alpha_{\mathrm{b}}}=b/d_{\mathrm{sep,0}}$ to specify the impact parameter. For the absolute value of the initial relative velocity, we assumed |${V}_{\mathrm{rel}}|=v_{200}$ in all cases.
Owing to the required computational time and substantial storage requirements, only a few dozen scenarios with more or less arbitrary parameter choices were feasible. Thus, our simulation suite covers some typical scenarios, such as edge-on (group 0) and face-on (group III) collisions of aligned disks, and a selection of intermediate cases with nonzero relative inclinations of the disks. In groups IV and V, we investigated scenarios where the secondary galaxy moves along an inbound trajectory that is inclined with respect to the disk plane of the primary (similar to satellites of the Milky Way, such as the Magellanic Clouds). If feasible, the simulations were run until the post-merger stage. In group 0, this was achieved for Sim. No. 1 to 5 within 2.52.52.5\;2.5Gyr, but not for Sim. No. 6. In the other groups, reductions of the fraction of the time step implied by the Courant-Friedrichs-Lewy criterion required terminating the simulations already after 1.51.51.5\;1.5Gyr in order to save computational time. However, all simulations captured at least the first interaction between the galaxies.